the present age of father is equal to the sum of ages of his 5 children. 12 years hence, the sum of ages of his children will be twice the age of their father. find the percentage of the father.
Answer by macston(5194)

F=father's age S=sum of children's ages
F=S
2(F+12)=S+60 Substitute for S
2(F+12)=F+60
2F+24=F+60 Subtract 24 from each side
2F=F+36 Subtract F from each side
F=36 ANSWER: Father's current age is 36
CHECK
In 12 years, sum of children's age = twice fathers (Sum of children currently=father)
S+60=2(F+12) Substitute for s
F+60=2(F+12)
36+60=2(36+12)
96=2(48)
96=96